# Topological properties of curved spacetime Su–Schrieffer–Heeger model

### v2

@@ IV Analytical expression for the exact zero energy eigenfunctions
From the wave-packet dynamics, it is apparent that the CST SSH model can also mimic horizon physics. The next question is, does this model show a topological phase transition? Note that in the $\sigma=0$ limit, this model is identical to the usual SSH model and shows a topological phase transition when $t_{1}=t_{2}$ between topologically trivial phase ($t_{1}>t_{2}$) and topologically non-trivial phase ($t_{2}>t_{1}$). The consequence of the topologically non-trivial phase is that for $t_{2}>t_{1}$, the Hamiltonian $H$ (4) under open-boundary conditions has two zero-energy states, and they are localized at the two edges. Now, the automatic question arises, what will happen to the fate of those edge states for $\sigma>0$, will such edge states survive? First, we perform a numerical check in Fig. 3. Figures in the upper panel suggest that zero energy states still survive for finite $\sigma$, though the gap is becoming smaller with increasing $\sigma$. Moreover, those states remain localized at the edges. However, unlike the usual SSH model, the states are asymmetric; the left edge state seems to become more localized compared to the right edge state with increasing $\sigma$.
It is reasonably straightforward to obtain those zero eigenstates by solving the time-independent Schrödinger equation for the Hamiltonian  (4), and one can easily obtain two recursive relations for two sub-lattices A and B.
@@
Another thing to note is the closing of the energy gap with $\sigma$. We observe that the gap gradually closes with increasing $\sigma$ and becomes almost non-existent as $\sigma$ reaches $1$. From Fig. 5, we observed that $\Delta E$ decreases with $\sigma$ as $\Delta E\propto N^{-\sigma}$ for a given fixed $t_{1}$ and $t_{2}$. Note that the finite-size system on the lattice will always have a gap, and a typical gap between two adjacent energy eigenvalues decreases with system size as $N^{-1}$ (we have found the same for our model as well). Thus, the ratio between $\Delta E$ and the typical gap in the spectrum scales as $N^{1-\sigma}$, indicating the presence of isolated zero-energy states even for non-zero $\sigma<1$. The fact that $N\Delta E$ increases with $N$, suggests the system remains gapped even in the thermodynamic limit for $\sigma<1$ [49]. Nevertheless, we also observe well-localized zero-energy edge states for the $\sigma$ higher than $1$ as well. This can be attributed to the robustness of the system’s topology.
Figure 4: for $t_{1}=0.5,t_{2}=1,$ ratio of right-to-left peak of probability densities vs $\sigma$ for system size $2N=1000$ ( 500 unit cells).Here the dotted lines represent the numerical calculations and the dashed lines are from the analytical expressions of the zero energy eigenstates 21,20
Figure 5: for $t_{1}=0.5,t_{2}=1$ energy gap $\Delta E$ at various system size NN for different $\sigma$
@@ V Topological signatures
In the previous section, the survival of the zero-energy edge states for $\sigma>0$ shows a signature that the CST-SSH Hamiltonian can also host the topologically non-trivial phase. In order to become more certain, one needs to understand the symmetries of the model, as well as some proper topological markers, which are required. First, we investigate whether even after the introduction of warping of spacetime $(\sigma)$, the Hamiltonian will still remain in the BDI category with winding number as its topological invariant quantity, as it is observed for the usual SSH model. Since the two position-dependent hopping terms $t_{1}(n)$ and $t_{2}(n)$ are real, the time reversal operator acting on the Hamiltonian $\hat{T}H\hat{T}^{-1}=H$ leaves it invariant. Moreover, the Hamiltonian being a spinless fermionic system, $\hat{T}^{2}=\mathbb{I}$. For chiral symmetry, first, we define $\hat{\Gamma}_{A}=\sum_{n=1}^{N}c_{n,A}^{\dagger}c_{n,A}$ and $\hat{\Gamma}_{B}=\sum_{n=1}^{N}c_{n,B}^{\dagger}c_{n,B}$ as projectors on sublattice A and B [50]. Then, the chiral operator is defined by $\hat{\Gamma}=\hat{\Gamma}_{A}-\hat{\Gamma}_{B}$, which anti-commutes with the Hamiltonian $\hat{\Gamma}H\hat{\Gamma}=-H$, since the bipartite nature is maintained in the modified position-dependent SSH model. This anti-commutation relation holds regardless of the position-dependent nature of the hopping amplitudes [50]. Time reversal and Chiral symmetry being present with $T^{2}=\Gamma^{2}=\mathbb{I}$, we can combine them as, $\hat{C}=\hat{T}\hat{\Gamma}$, which can acts as a particle hole symmetry, i.e., an anti-unitary operator that anti-commutes with the Hamiltonian, and also $\hat{C}^{2}=\mathbb{I}$. Thus, this CST-SSH Hamiltonian remains in the BDI category in the Altland-Zirnbauer tenfold classification of topological insulators and superconductors [51, 52] with an integer $\mathbb{Z}$ topological invariant like its standard counterpart in the usual SSH model. So, the symmetries remain invariant under the position-dependent hopping amplitude, and so does the topological invariant, i.e., winding number. However, in the CST-SSH model, the absence of translational symmetry prevents us from utilizing the k-space to calculate the winding number. Therefore, we adopt two real space markers to calculate the topological invariants: (i) Local topological marker (LTM), (ii) Mean chiral displacement (MCD).
@@ V.1 Local Topological Marker
Figure 6: Local Topological Marker (LTM) for $(a)\sigma=0$ and, $(b)\sigma=1$ for trivial $(t_{1},t_{2})=(1,0.5)$ and topological phase $(t_{1},t_{2})=(0.5,1)$, for the CST-SSH model having system size $2N=1000$ (500 unit cells)
Figure 7: (a).(b): Time dependent MCD for both topological (red) with $(t_{1},t_{2})=(0.5,1)$ and trivial (blue) with $(t_{1},t_{2})=(1,0.5)$for (a) $\sigma=0$ and,(b) $\sigma=1$ for the CST-SSH model having system size $2N=1000$ (500 unit cells) , (c)Fluctuation measurement as Var($C(t)$) vs $\sigma$ for different system sizes NN
@@
Here, $\mathbb{X}$ is the position operator. $\mathbb{Q}$ can be defined from the modal matrix $\mathbb{U}$, which is comprised of all the normalized eigen vectors with ascending order, explicitly, $\mathbb{U}=[\mathbf{U}_{1},\mathbf{U}_{2},...\mathbf{U}_{n},\mathbf{U}_{n+1},...\mathbf{U}_{N}]$. Here, $\mathbb{U}_{-}=[\mathbf{U}_{1},\mathbf{U}_{2},...\mathbf{U}_{n}],$ and $\mathbb{U}_{+}-=[\mathbf{U}_{n+1},\mathbf{U}_{n+2},...\mathbf{U}_{N}]$ are corresponds to below and above the band gap energy spectrum. Then one can define the projectors as $\mathbb{P}_{-}=\mathbb{U}_{-}\mathbb{U}_{-}^{T}$ and $\mathbb{P}_{+}=\mathbb{U}_{+}\mathbb{U}_{+}^{T}$, and $\mathbb{Q}$ as, $\mathbb{Q}=\mathbb{P}_{+}$-$\mathbb{P}_{-}$, which further can be decomposed as, $\mathbb{Q}=\mathbb{Q}_{AB}+\mathbb{Q}_{BA}=\Gamma_{A}\mathbb{Q}\Gamma_{B}+\Gamma_{B}\mathbb{Q}\Gamma_{A}$, where $\Gamma=\Gamma_{A}-\Gamma_{B}$ is the chiral operator. More about the formula and the operators can be found in the appendix of [53, 54]. From the figure 6, we observe that even for $\sigma=1$, the LTM shows the same winding number for the CST-SSH Hamiltonian as its usual SSH counterpart. While $t_{2}>t_{1}$, the LTM shows $\nu=1$, indicating topologically non-trivial phase, and for $t_{1}>t_{2}$, $\nu=0$ signifying the trivial phase. There is no apparent difference between the standard and the position-dependent hopping SSH model for this static marker.
@@ V.2 Mean Chiral Displacement
The previous topological measure was a static measure; in this section, we use a dynamical topological measure, which we call Mean Chiral Displacement (MCD) [55, 53], can be used to detect the winding number. The MCD is defined as ,
@@ VI Quench dynamics: from topological to trivial phase
[…]
Figure 9: For the CST-SSH model having system size $2N=1000$ (500 unit cells), Time evolution of the zero energy eigenstates: (a) left, and, (b) right edge zero energy eigenstate of a Hamiltonian with winding number 1, in another Hamiltonian with winding number zero. Inset of Fig. (a) shows the semiclassical dynamics near the origin, where the left edge state is located. Inset of Fig. (b) shows the semiclassical dynamics where the right edge state is located.
In this section, we study the quench dynamics across the topological transition point. We prepare our initial state $|\psi_{i}\rangle$ as one of the zero-energy eigenstates of the CST-SSH Hamiltonian for $t_{2}>t_{1}$. Note, this is the topologically non-trivial phase, having two edge modes. Then, we quench across the transition point in the topologically trivial phase. It implies $t_{2}<t_{1}$ for the post-quench Hamiltonian, which we identify as $H_{f}$. The survival probability of the initial state after quench is given by [56, 57],
@@
It is basically a measurement of the likelihood of the zero-energy eigenstate of a topologically nontrivial Hamiltonian remaining the same after the unitary time evolution of the state in another Hamiltonian with winding number 0. Figure. 8 shows the variation of the survival probability with time for the usual SSH model, i.e, $\sigma=0$, and for the CST-SSH model with $\sigma=0.5$. We quench from $(t_{1},t_{2})=(0.5,1)$ to $(t_{1},t_{2})=(1,0.5)$. For the initial state, i.e., localized at the right edge, the time evolution of the survival probability is almost identical for the usual SSH and the CST-SSH model (see Fig. 8a). The survival probability goes to zero reasonably quickly. However, the dynamics are quite different between these two models for the left edge mode (see Fig. 8b). While for the usual SSH, the dynamic is almost indistinguishable from that obtained for the right edge mode, for $\sigma>0$, the survival probability seems to oscillate between 0 and $1$ for the CST-SSH model. We find that the oscillation period increases with increasing $\sigma$. The fact that dynamics is identical for usual SSH and not for CST-SSH is not too surprising, given that we found in Sec. IV that both edge states are symmetric for $\sigma=0$, and asymmetry kicks in as soon as $\sigma>0$. However, it still does not explain why the survival probability of the left-edge states oscillates. Hence, we plot the absolute value of the time-evolved wavefunction in Fig. 9, and find that for the CST-SSH model, while under time evolution, the right edge state moves toward the other side of the lattice, while the left edge mode kind of remains there. If one tracks the peak position, one finds that it moves a bit to the right and then again comes back, keeping up this to-and-fro motion. This is precisely what gets manifested in the survival probability plot. Moreover, we find the same picture even when solving the semi-classical equation of motion (see inset of Fig. 9).
VII Extended CST-SSH Model
[…]
In this section, we consider an extended version of the SSH model, described by the Hamiltonian 5. Here, this system can possess a winding number ranging from 1-1 to $2$. Such models can host more than one edge mode on each lattice edge (see Appendix A). While in the case of the CST-SSH model, the winding numbers of the distinct topological phases were either $\nu=0$ (topologically trivial) or $\nu=1$. By choosing suitable parameters, one can have even topological phases with a higher winding number in case of the Hamiltonian extended CST-SSH model in Eq.5. First, we consider $t_{3}=t_{4}=0$, and in this limit for the Hamiltonian $H_{ext}$, we have also calculated the local topological marker (LTM) and find in case of both $\sigma=0$ and $1$, the winding numbers are $\nu=0$ and $1$ for $(t_{1},~t_{2})~=~(1,~0.5),~(0.5,~1)$, respectively. Likewise, for parameters $(t_{1},~t_{2},~t_{3},~t_{4})~=~(0.5,~0,~1,~0)$, (0,1,1,3)~(0,1,1,-3), one obtains winding number $\nu=-1,~2$, respectively for both $\sigma=0$ and $\sigma=1$ (see Fig.  10). These results show that topological features in the CST version of the SSH type models survive even for the extended CST-SSH model. Thus, we can say that the two gap-closing constraints also remain the same for the extended CST-SSH model.
In this extended CST-SSH model, we can see clear winding number transition points. For example, $(t_{1},t_{2},t_{3},t_{4})=(3,-2,1,-4)$ is a transition point between $\nu=1$ to $\nu=2$ phase provided we keep $t_{1}$, $t_{2}$, $t_{3}$ fixed and vary $t_{4}$. If we allow $t_{4}$ to go slightly below 4-4, while keeping the other parameters the same, the winding number becomes $2$, and if one allows $t_{4}$ to go higher than 2-2, it will make the winding number $1$; making the particular combination of hoppings a transition point. If we study a Gaussian wavepacket evolution in this extended CST-SSH Hamiltonian for this particular topological transition point, we observe, in a similar manner, an eternal slowdown of the wavepacket for $\sigma>0$, for initial wavepacket momentum $|p_{0}|=\frac{\pi}{2}$ (see Fig.11a). Figure. 11b also shows that for another transition point obeying $t_{1}+t_{2}+t_{3}+t_{4}=0$ with $(t_{1},t_{2},t_{3},t_{4})=(0,1,1,-2)$, also shows critical slowdown for $|p_{0}|=\pi$ . For other than transition points, the wavepacket returns before reaching the origin (see Fig.11c). Similar behavior can be observed for other transition points, like from $\nu=0$ to $\nu=1$, from $\nu=0$ to $\nu=-1$, from $\nu=0$ to $\nu=2$ etc. Thus, these observations imply the presence of a synthetic horizon at the topological phase-transition points, and that strengthens our claims that horizon physics emerges only at the topological transition points in these models.
VII.1 Semiclassical dynamics for the Extended CST-SSH model
In a similar spirit to the previous section. III, we can analyze the semiclassical dynamics for the extended version of the CST-SSH model. Since here we have two additional hoppings, in our recurring relation, we will have two more variables. The two relations will be:
$=$ $\epsilon\psi_{A,n}$
Now, we, in a similar manner, after introducing the continuous function $\tilde{\psi}_{A/B}(x_{n})$, and expanding it in Taylor series, can obtain the continuum Hamiltonian, and then similarly the semiclassical equations of motion:
$\tilde{H}_{ext}=\begin{pmatrix}0&T(\hat{x},\hat{p})\\[2.84526pt] T^{*}(\hat{x},\hat{p})&0\end{pmatrix}$,
[…]
Here, The energy expression,
$E_{\pm}(x,p)=\pm|T(x,p)|=\pm\sqrt{T_{R}^{2}(x,p)+T_{I}^{2}(x,p)}$
, where, $T_{R}=t_{1}(x)+(t_{2}(x)+t_{3}(x))\text{cos}(2p)+t_{4}(x)\text{cos}(4p)$ and $T_{I}(x,p)=(t_{3}(x)-t_{2}(x))\text{sin}(2p)-t_{4}(x)\text{sin}(4p)$ are the real and imaginary part of $T(x,p)$. Now, after substituting the values of $T_{R}$ and $T_{I}$, one can obtain the equations of motion as:
where, $\partial_{p}$ and $\partial_{x}$ are the partial derivatives with respect to $p$ and $x$, respectively. Fig. 11d shows the comparison between the numerical time evolution ( dotted black lines) and the semiclassical equations of motion ( colored lines ) to obtain the peak positions of the wave packets at various times for different hopping parameters. The semiclassical calculations show excellent agreement with the numerical time evolution for this extended version as well. One can obtain the analytical value of the turning point following a similar way, by equating the energy expression at $(x_{0},p_{0})$ and at $(x_{\min},p)$, where choosing an appropriate $p$ makes $\dot{x}=0$. For $p=0$, the turning position reads as:
$x_{\min}=x_{0}\Biggl{[}\frac{(t_{1}+(t_{2}+t_{3})\cos 2p+t_{4}\cos 4p)^{2}+((t_{3}-t_{2})\sin 2p-t_{4}\sin 4p)^{2}}{(t_{1}+t_{2}+t_{3}+t_{4})^{2}}\Biggr{]}^{\frac{1}{2\sigma}}$ (30)
This gives us that, when $t_{1}-(t_{2}+t_{3})+t_{4}=0$, $x_{\min}$ is equal to 0 at $|p_{0}|=\frac{\pi}{2}$; and when $(t_{1}-t_{4})^{2}+(t_{3}-t_{2})^{2}=0$, $x_{\min}=0$ at $|p_{0}|=\frac{\pi}{4}$, which are precisely two of the gap closing conditions, i.e., topological transition constraints. One important thing to observe is that when we derived the formula, we already assumed that $(t_{1}+t_{2}+t_{3}+t_{4})\neq 0$, which prevents us from obtaining the other condition. So, if our Hamiltonian has the hopping parameters $t_{1}$, $t_{2}$, $t_{3}$, $t_{4}$ that add to 0, we can not use the formula. In such cases, we go to the next root of $\sin(2p)$, which gives us $\dot{x}=0$, which is for $p=\pi/2$. Then the $x_{\min}$ will have $(t_{1}-(t_{2}+t_{3})+t_{4})^{2}$ in the denominator instead. In this formula, $x_{\min}$ will be 0 only when $(t_{1}+t_{2}+t_{3}+t_{4})=0$, at initial momentum $|p_{0}|=\pi$, which is the other gap closing condition. In case of non-transition points, the minimum of these two would give us the turning point.
Thus, we see that the gap-closing conditions that determine the topological transition between different winding numbers are also conditions for $x_{\min}$ to become 0. This shows that only at topological transition points, we observe horizon physics at the extended CST-SSH model, which strengthens our findings, discussed in the above sections.

### v3

@@ IV Analytical expression for the exact zero energy eigenfunctions
From the wave-packet dynamics, it is apparent that the extended CST-SSH model can also mimic horizon physics for certain choices of parameters. This naturally raises the question of whether this model also exhibits a topological phase transition. Note that in the $\sigma=0$ limit, this model reduces to the usual extended SSH model, which undergoes a topological phase transition between a topologically trivial phase and a topologically non-trivial phase. In the topologically non-trivial phase, under open-boundary conditions, zero-energy states appear that are localized at the two edges of the system. An immediate question then arises: what happens to these edge states when $\sigma>0$? In particular, do such edge states survive in the presence of the CST deformation? In the following, we analyze this question. For analytical simplicity, we restrict our discussion to the CST-SSH model; however, the analysis can be straightforwardly extended to the extended CST-SSH model.
First, we perform a numerical check in Fig. 3. Figures in the lower panel suggest that zero energy states still survive for finite $\sigma$, though the upper panel figures showing gap is becoming smaller with increasing $\sigma$. Moreover, those states remain localized at the edges. However, unlike the usual SSH model, the states are asymmetric; the left edge state seems to become more localized compared to the right edge state with increasing $\sigma$.
It is reasonably straightforward to obtain those zero eigenstates by solving the time-independent Schrödinger equation for the CST-SSH Hamiltonian, and one can easily obtain two recursive relations for two sub-lattices A and B.
@@
Another thing to note is the closing of the energy gap ($\Delta E$) with $\sigma$. We observe that the gap gradually closes with increasing $\sigma$ and becomes almost non-existent as $\sigma$ reaches $1$. Such gaplessness in topological insulator due to inhomogeneity or disorder has also been observed in the literature  [64, 62, 37]. From Fig. 5, we observed that $\Delta E$ decreases with $\sigma$ as $\Delta E\propto N^{-\sigma}$ for a fixed $t_{1},t_{2},t_{3},t_{4}$. Note that the finite-size system on the lattice will always have a gap, and a typical gap between two adjacent energy eigenvalues decreases with system size as $N^{-1}$ (we have found the same for our model as well). Thus, the ratio between $\Delta E$ and the typical gap in the spectrum scales as $N^{1-\sigma}$, indicating the presence of isolated zero-energy states even for non-zero $\sigma<1$. The fact that $N\Delta E$ increases with $N$, suggests the system remains gapped even in the thermodynamic limit for $\sigma<1$ [36], and gapless for $\sigma\geq 1$ (see Fig. 5 and also Appendix A). This observation indicates that for any combination of $t_{1},t_{2},t_{3},t_{4}$, even if the parameters do not correspond to a topological phase transition point, the system becomes gapless in the $\sigma\geq 1$ region. However, in this regime we do not observe horizon physics. Hence, the presence of a gapless spectrum alone does not guarantee the occurrence of a critical slowdown. Furthermore, unlike the standard SSH model, the extended SSH model cannot, at the topological transition points, be mapped onto a single-band tight-binding CST Hamiltonian. Nevertheless, we also observe well-localized zero-energy edge states for the $\sigma$ higher than $1$ as well.
Figure 4: Ratio of right-to-left peak of probability densities vs $\sigma$ for a system size $2N=1000$ (500 unit cells) with $(t_{1},t_{2},t_{3},t_{4})=(0.5,1,0,0)$. Here the dotted lines represent the numerical calculations and the dashed lines are from the analytical expressions of the zero energy eigenstates 21,20
Figure 5: Energy gap $\Delta E$ at various system size NN for different $\sigma$, for the Extended CST-SSH with $(t_{1},t_{2},t_{3},t_{4})=(2,1,-1,-4)$ (top), and, for $(t_{1},t_{2},t_{3},t_{4})=(0.5,1,0,0)$ ( bottom)
@@ V Topological signatures
In the previous section, the survival of the zero-energy edge states for $\sigma>0$ shows the signature that the CST version of the SSH model can retain the imprint of the same topological phases and transition of the usual SSH models. However, the CST-SSH models remain gapless for $\sigma\geq 1$. In order to understand that further, one needs to understand the symmetries of the model, as well as some proper topological markers, which are required. First, we investigate whether even after the introduction of warping of spacetime $(\sigma)$, the Hamiltonian will still remain in the BDI category with winding number as its topological invariant quantity, as it is observed for the usual SSH model. Since the position-dependent hopping terms are real, the time reversal operator acting on the Hamiltonian $\hat{T}H\hat{T}^{-1}=H$ leaves it invariant. Moreover, the Hamiltonian being a spinless fermionic system, $\hat{T}^{2}=\mathbb{I}$. For chiral symmetry, first, we define $\hat{\Gamma}_{A}=\sum_{n=1}^{N}c_{n,A}^{\dagger}c_{n,A}$ and $\hat{\Gamma}_{B}=\sum_{n=1}^{N}c_{n,B}^{\dagger}c_{n,B}$ as projectors on sublattice A and B [2]. Then, the chiral operator is defined by $\hat{\Gamma}=\hat{\Gamma}_{A}-\hat{\Gamma}_{B}$, which anti-commutes with the Hamiltonian $\hat{\Gamma}H\hat{\Gamma}=-H$, since the bipartite nature is maintained in the modified position-dependent SSH models. This anti-commutation relation holds regardless of the position-dependent nature of the hopping amplitudes [2]. Time reversal and Chiral symmetry being present with $T^{2}=\Gamma^{2}=\mathbb{I}$, we can combine them as, $\hat{C}=\hat{T}\hat{\Gamma}$, which can acts as a particle hole symmetry, i.e., an anti-unitary operator that anti-commutes with the Hamiltonian, and also $\hat{C}^{2}=\mathbb{I}$. Thus, the CST-SSH Hamiltonians remain in the BDI category in the Altland-Zirnbauer tenfold classification of topological insulators and superconductors [1, 46] with an integer $\mathbb{Z}$ topological invariant like its standard counterpart in the usual SSH model. Therefore, the symmetries remain invariant under the position-dependent hopping amplitude, and so does the topological invariant, i.e., winding number. However, in the CST-SSH models, the absence of translational symmetry prevents us from utilizing the k-space to calculate the winding number. Therefore, we adopt two real space markers to calculate the topological invariants: (i) Local topological marker (LTM), (ii) Mean chiral displacement (MCD).
@@ V.1 Local Topological Marker
Figure 7: Local Topological Marker (LTM) for $(a)\sigma=0$ and, $(b)\sigma=1$ for trivial $(t_{1},t_{2},t_{3},t_{4})=(1,0.5,0,0)$ and topological phase $(t_{1},t_{2},t_{3},t_{4})=(0.5,1,0,0)$, for the CST-SSH model having system size $2N=1000$ (500 unit cells)
Figure 8: Time dependent MCD for both topological (red) with $(t_{1},t_{2},t_{3},t_{4})=(0.5,1,0,0)$ and trivial (blue) with $(t_{1},t_{2},t_{3},t_{4})=(1,0.5,0,0)$ for (a) $\sigma=0$ and,(b) $\sigma=1$ for system size $2N=1000$ (500 unit cells)
Figure 9: Measurement of survival probability $P_{i}(t)$ from winding number 1 (with $(t_{1},t_{2},t_{3},t_{4})=(0.5,1,0,0)$) to winding number 0 (with $(t_{1},t_{2},t_{3},t_{4})=(1,0.5,0,0)$) for (a) right edge zero energy eigen state and, (b) left edge zero energy eigen state, for system size $2N=1000$ (500 unit cells)
Figure 10: Time evolution of the zero energy eigenstates for system size $2N=1000$ (500 unit cells), : (a) left, and, (b) right edge zero energy eigenstate of a Hamiltonian with winding number 1, in another Hamiltonian with winding number zero. Inset of Fig. (a) shows the semiclassical dynamics near the origin, where the left edge state is located. Inset of Fig. (b) shows the semiclassical dynamics where the right edge state is located.
Figure 11: (a)Values of slowdown time scale $\tau_{s}$ with $t_{1}=1,t_{2}=0.5$ and over range of $t_{3},t_{4}$ . Here $\nu_{i},(i=-1,0,1,2)$ represent regions with winding number $i$. (b) for a higher order gap closing point (1,0.5,-0.5,-1), two different critical slowdown parameters $\tau_{s}$ for two initial critical momenta, inset plot showing a Log-linear plot of the same, with straight-line fits used to extract $\tau_{s}$ for each $p_{0}$, For system size $2N=1000$ (500 unit cells):
@@
Here, $\mathbb{X}$ is the position operator. $\mathbb{Q}$ can be defined from the modal matrix $\mathbb{U}$, which is comprised of all the normalized eigen vectors with ascending order, explicitly, $\mathbb{U}=[\mathbf{U}_{1},\mathbf{U}_{2},...\mathbf{U}_{n},\mathbf{U}_{n+1},...\mathbf{U}_{N}]$. Here, $\mathbb{U}_{-}=[\mathbf{U}_{1},\mathbf{U}_{2},...\mathbf{U}_{n}],$ and $\mathbb{U}_{+}-=[\mathbf{U}_{n+1},\mathbf{U}_{n+2},...\mathbf{U}_{N}]$ are corresponds to below and above the band gap energy spectrum. Then one can define the projectors as $\mathbb{P}_{-}=\mathbb{U}_{-}\mathbb{U}_{-}^{T}$ and $\mathbb{P}_{+}=\mathbb{U}_{+}\mathbb{U}_{+}^{T}$, and $\mathbb{Q}$ as, $\mathbb{Q}=\mathbb{P}_{+}$-$\mathbb{P}_{-}$, which further can be decomposed as, $\mathbb{Q}=\mathbb{Q}_{AB}+\mathbb{Q}_{BA}=\Gamma_{A}\mathbb{Q}\Gamma_{B}+\Gamma_{B}\mathbb{Q}\Gamma_{A}$, where $\Gamma=\Gamma_{A}-\Gamma_{B}$ is the chiral operator. More about the formula and the operators can be found in the appendix of [50, 33]. From the Figs. 6, 7, we observe that even for $\sigma=1$, the LTM shows the same winding number for the both CST-SSH Hamiltonians as its usual SSH counterparts, for the exact values of the hopping. There is no apparent difference between the standard and the position-dependent hopping SSH models for this static marker, also indicating same topological phase boundaries for the CST models as well.
@@ V.2 Mean Chiral Displacement
The previous topological measure was a static measure; in this section, we use a dynamical topological measure, which we call Mean Chiral Displacement (MCD) [8, 50], can be used to detect the winding number which is defined as,
@@ VI Quench dynamics: from topological to trivial phase
In this section, we study the quench dynamics across the topological transition point. For the sake of numerical simplicity and understading the mechanism better, we study the dynamics in the limit of $t_{3}=t_{4}=0$, leaving the Hamiltonian with two possible winding numbers: 0,1. We prepare our initial state $|\psi_{i}\rangle$ as one of the zero-energy eigenstates of the CST-SSH Hamiltonian for $t_{2}>t_{1}$. Note, this is the topologically non-trivial phase, having two edge modes. Then, we quench across the transition point in the topologically trivial phase. It implies $t_{2}<t_{1}$ for the post-quench Hamiltonian, which we identify as $H_{f}$. The survival probability of the initial state after quench is given by [18, 52],
@@
It is basically a measurement of the likelihood of the zero-energy eigenstate of a topologically nontrivial Hamiltonian remaining the same after the unitary time evolution of the state in another Hamiltonian with winding number 0. Although the topological markers produce nearly identical results for the CST-SSH and the conventional SSH models, we had identified a clear asymmetry in the edge modes of the CST-SSH model, unlike the identical edge modes of the usual SSH model. In this section, our goal is to examine how these asymmetric edge modes evolve under our quench protocol. By contrast, in the usual SSH model the symmetric edge modes are expected to exhibit identical dynamics under the same protocol.
@@
Figure. 9 shows the variation of the survival probability with time for the usual SSH model, i.e, $\sigma=0$, and for the CST-SSH model with $\sigma=0.5$. We quench from $(t_{1},t_{2})=(0.5,1)$ to $(t_{1},t_{2})=(1,0.5)$. For the initial state, i.e., localized at the right edge, the time evolution of the survival probability is almost identical for the usual SSH and the CST-SSH model (see Fig. 9a). The survival probability goes to zero reasonably quickly. However, the dynamics are quite different between these two models for the left edge mode (see Fig. 9b). While for the usual SSH, the dynamic is almost indistinguishable from that obtained for the right edge mode, for $\sigma>0$, the survival probability seems to oscillate between 0 and $1$ for the CST-SSH model. We find that the oscillation period increases with increasing $\sigma$. The fact that dynamics is identical for usual SSH and not for CST-SSH is not too surprising, given that we found in Sec. IV that both edge states are symmetric for $\sigma=0$, and asymmetry kicks in as soon as $\sigma>0$. However, it still does not explain why the survival probability of the left-edge states oscillates. Hence, we plot the absolute value of the time-evolved wavefunction in Fig. 10, and find that for the CST-SSH model, while under time evolution, the right edge state moves toward the other side of the lattice, while the left edge mode kind of remains there. If one tracks the peak position, one finds that it moves a bit to the right and then again comes back, keeping up this to-and-fro motion. This is precisely what gets manifested in the survival probability plot. Moreover, we find the same picture even when solving the semi-classical equation of motion (see inset of Fig. 10).
VII Quantification of the critical slowdown
At this point, we have established that there exists a connection between the topological transition points in SSH models and the observation of horizon physics in their CST counterparts. Unlike the standard SSH model, the extended SSH model supports higher-order topological phases. While obtaining the semiclassical trajectories of the wave packets, we observed that at different topological phase transition points the rate at which the wave packet approaches the origin varies across the hopping parameter space. Therefore, in this section we quantify the slowdown rate by introducing a characteristic time scale, $\tau_{s}$, which allows us to better understand the horizon physics in topological systems with multiple gap-closing constraints for usual extended SSH models. During the critical slowdown, the peak position of the wave packet follows
$x(\tau)=x_{0}e^{-\tau/\tau_{s}},$ (25)
as shown in Fig. 11b. Thus, for a chosen range of initial hopping parameters, we can analyze how fast or slow the wave packet moves toward the origin. Figure 11a shows the different winding numbers and the transition lines separating the topological phases. The heatmap indicates how slowly the wave packet moves toward the origin: larger values of $\tau_{s}$ correspond to slower motion. We observe that the horizon physics, characterized by the critical slowdown, appears only along the transition lines.
In our calculations we fix $t_{1}=1$ and $t_{2}=0.5$, and compute the slowdown parameter $\tau_{s}$ over a range of $t_{3}$ and $t_{4}$. We find that when condition (iii) and either condition (i) or (ii) of Eq. 5 are satisfied simultaneously, the wave packet does not evolve with time and effectively becomes stationary. The reason is that satisfying two gap-closing conditions at the same Dirac point makes the initial group velocity vanish. In the vicinity of these stationary points, the value of $\tau_{s}$ becomes larger, implying that the wave packets approach the origin more slowly compared to other transition points where only a single gap-closing constraint is satisfied. This behavior is visible in Fig. 11a as the bright yellow regions near the junctions. We also observe another interesting phenomenon. At higher-order gap-closing points of the usual extended SSH model where both conditions (i) and (ii) of Eq. 5 are satisfied, two different critical slowdowns emerge for two different initial momenta $p_{0}$, resulting in two distinct slowdown parameters $\tau_{s}$ (see Fig. 11b). Such features arise only in topological systems with higher-order gap closings, such as the extended CST-SSH model. Note that we would like to clarify that whenever we refer to gap-closing points in this section, we mean the gap-closing points of the usual extended SSH model at the topological transition points, and not those of the CST-SSH model. This distinction is important because the CST-SSH model remains gapless for $\sigma\geq 1$, as discussed previously. Finally, this provides an additional degree of freedom to control the critical slowdown in experimental setups designed to mimic black hole horizon physics. Such flexibility is absent in the simpler CST-SSH model.